OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -8, an = 5 • an-1 @agent0smith

OpenStudy (agent0smith):

Do this like this one http://openstudy.com/study#/updates/52d31139e4b01e5fc1dfd19e

OpenStudy (agent0smith):

Plug in a1 to find a2. Then plug in a2 to find a3, and so on.

OpenStudy (anonymous):

Don't hate me but can we do it together because honestly like not understanding or an example

OpenStudy (agent0smith):

How'd you do that last one correctly? You can do this in the same way.

OpenStudy (anonymous):

someone sorta just gave me the answer

OpenStudy (agent0smith):

\[\Large a_1 = -8, \text{ and } a_n = 5 • a_{n-1}\]Find a2: \[\Large a_2 = 5 • a_{2-1} = 5•a_1\]and you know what a1 is.

OpenStudy (agent0smith):

Once you have a2, find a3:\[\Large a_3 = 5 • a_{3-1} = 5 • a_{2}\]

OpenStudy (anonymous):

i got: -8, -40, -200, -1000, -5000, -25,000

OpenStudy (agent0smith):

Good, as long as you worked them out yourself this time and weren't given them :P

OpenStudy (anonymous):

I did i basically just get multiplying the number before by 5

OpenStudy (agent0smith):

Exactly! :)

OpenStudy (agent0smith):

an is the term you want. an-1 is just the term before it.

OpenStudy (anonymous):

Ok thanks c: